OpenStudy (anonymous):

In a table, how do I tell if it's linear, expodential, Inversly or dirrectly?

OpenStudy (anonymous):



OpenStudy (science0229):

First, if this is varying directly or indirectly, the table would have to follow y= kx or y=k/x (k=constant)

OpenStudy (science0229):

k cannot be 0

OpenStudy (science0229):

There is no values of k that satisfy the first point. k/1=0 -----> k=0 (But we said that k cannot be 0) So, this is not varying directly

OpenStudy (anonymous):

is that the same thing as mx as in: y=mx+b

OpenStudy (science0229):

Similar proof for why it can't be direct variation.

OpenStudy (anonymous):

Ok..... So how do I figure out the rest?

OpenStudy (science0229):

Linear form is as you said: y=mx+b

OpenStudy (science0229):

If this follows the linear form, all points have to be on that equation.

OpenStudy (science0229):

first point: m+b=0 second point: 2m+b=2 third point: 3m+b=4

OpenStudy (anonymous):

is there a way to look at the table and immediatly know which one it is

OpenStudy (science0229):

and so on